# Spin transport and lack of quantisation in the $A\mathrm{II}$ class on the honeycomb structure

### v1

@@ Abstract.
We investigate spin transport in a class of two-dimensional $A\mathrm{II}$ insulators on the honeycomb structure, the Kane–Mele model being an emblematic example in this class. We derive the spin conductivity by the linear response à la Kubo and show that it is well-defined and independent of the choice of the spin current. For models that do not conserve the spin, we demonstrate that the deviation of the spin conductivity from the quantised value is, at worst, quadratic in the spin-non-conserving terms, thus improving previous results. Additionally, we show that the leading-order corrections are actually quadratic for some models in the class, demonstrating that the spin conductivity is not universally quantised. Consequently, our results show that, in general, there is no direct connection between the spin conductivity and the Fu–Kane–Mele index.
@@ 1. Introduction
For spin-conserving systems, the spin Hall conductivity is connected with the Fu–Kane–Mele index, as is the case for the charge Hall conductivity with the Chern number: in particular, the spin Hall conductivity is quantised in units of $e/h$, the quantisation integer modulo two being the Fu–Kane–Mele index [BZ, HK]. However, in spin-non-conserving systems, it remains both experimentally and theoretically unclear whether any spin response function can be associated with the Fu–Kane–Mele index. On the one hand, experimental measurements are imprecise, likely because of the presence of magnetic impurities that partially break time-reversal symmetry [Mol1, Mol2]. On the other hand, theoretical problems already arise in identifying a suitable spin-current operator, the system lacking the associated conserved quantity [SZXN]. In this regard, two operators serve as natural generalizations of the spin current for spin-conserving systems: the conventional spin current, defined as the product of the charge current with the spin operator, and the proper spin current introduced in [SZXN, ZWSXN], which is tied with a continuity equation and with the Onsager relations.
@@
In this work, using the Kubo formula, we continue the study of the spin conductivity initiated in [MaPaTa, MaPaTe, MaMo], by specialising to a class of models on the two-dimensional honeycomb structure which we denote by $\mathrm{AII}^{\hexagon}$. This class of models consists of the one-body Hamiltonians in the Altland–Cartan–Zirnbauern $\mathrm{AII}$ class [AZ, RSFL], additionally characterised by the spatial symmetries of the Kane–Mele model [KM1, KM2]. For a precise definition, see Definition 2.4. The presence of spatial symmetries is crucial in our analysis to remove the ill-posedness problems and the ambiguities related to the choice of the spin current. Unlike [MaPaTe], in which the NEASS approach is used, here we follow Kubo’s approach to transport coefficients. Moreover, we exploit the spatial symmetries in $\mathrm{AII}^{\hexagon}$ to prove that the spin conductivity is an antisymmetric tensor. This property makes the spin conductivity invariant under rotations, and thus establishes that it measures an intrinsic transverse response, independent of the orientation of the laboratory [SVWBJ].
In addition to reviewing key well-posedness findings, we prove two novel results for insulators in the $\mathrm{AII}^{\hexagon}$ class that almost conserve the spin. First, we show that the deviation of the spin conductivity from the quantised value (which is observed in spin-conserving systems) is, at worst, quadratic in the spin-non-conserving terms, improving previous results [Sch, MaPaTe]. Note that in theoretical units, which we henceforth adopt, quantisation of the spin conductivity means valued in $\frac{1}{2\pi}\mathbb{Z}$. Additionally, we show that in certain models the leading-order corrections to the quantised value are indeed quadratic, implying that spin conductivity is not universally quantised. As a result, our findings demonstrate that no general connection exists between spin conductivity and the Fu–Kane–Mele index when implementing the linear response by Kubo’s formula. To the best of our knowledge, this is the first rigorous result in this direction.
Let us now state these results more precisely. We refer to insulators in the $\mathrm{AII}^{\hexagon}$ class as pairs $(H,\mu)$ where $H$ is a Hamiltonian in $\mathrm{AII}^{\hexagon}$ with a spectral gap, and where $\mu\in\mathbb{R}$ is in the spectral gap. Let us denote by $S_{z}=\frac{1}{2}\sigma_{z}$ the spin operator in the $z$ direction. To quantify spin conservation, we follow [Sch, Eq. (1)] and split any Hamilton operator $H$ as follows
@@
    For any short-range insulator $(H,\mu)$ in the $\mathrm{AII}^{\hexagon}$ class, the spin conductivity tensor, which we denote by $\sigma^{s}(H,\mu)=\big{(}\sigma^{s}_{ij}(H,\mu)\big{)}_{i,j=1,2}$, is well-defined and independent of the choice of the spin current operator. Furthermore, it holds true that $\sigma^{s}_{ij}(H,\mu)=-\sigma^{s}_{ji}(H,\mu)$, for any $i,j=1,2$.
@@
    For any short-range insulator $(H,\mu)$ in the $\mathrm{AII}^{\hexagon}$ class that almost conserves the spin, there exists a constant $C$ independent of ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$ such that
    $\mathrm{dist}\Big{(}\sigma^{s}_{12}(H,\mu)\,,\,\frac{1}{2\pi}\mathbb{Z}\Big{)}\leq C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}.$
@@
    There exist insulators $(H,\mu)$ in the $\mathrm{AII}^{\hexagon}$ class that almost conserve the spin, and a constant $C>0$ independent of ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$ such that
    $\mathrm{dist}\Big{(}\sigma^{s}_{12}(H,\mu)\,,\,\frac{1}{2\pi}\mathbb{Z}\Big{)}\geq C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|[H,S_{z}]\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}^{2}.$
Each item of Theorem 1.1 is elaborated in different sections of the manuscript: part $(i)$ is addressed in Propositions 3.13 and 3.15, part $(ii)$ in Theorem 3.22, and finally part $(iii)$ in Theorem 4.10. As previously mentioned, the proof of $(i)$ follows the ideas of [MaPaTa, MaPaTe], but relies on Kubo’s approach to the computation of transport coefficients. Our analysis shows that the derivation of the spin conductivity by Kubo’s formula agrees with the one obtained via NEASS, see Remark 3.18. The proof of $(ii)$ involves more refined algebraic manipulations than in previous work, allowing us to extract commutators between operators that nearly commute with the spin operator, thus providing improved estimates. However, these manipulations can only be applied once, meaning $(ii)$ does not rule out the possibility that the corrections could be universally of higher order. This is however not the case due to $(iii)$. To prove $(iii)$, we explicitly construct an extension of the Kane–Mele model with next-to-nearest-neighbour Rashba interaction (see (4.1)) and analyse the phase diagram for suitable parameter choices. In Proposition 4.9, we show that, like in the Kane–Mele model, this diagram features three insulating phases separated by two semi-metallic phases (i.e., a metallic phase with a point-like Fermi surface). Using the “imaginary-time” representation of the Kubo formula for the spin conductivity, established via a Wick rotation (as in [GJMP, GMP2]) and detailed in Theorem 4.15, we compute the discontinuity of the spin conductivity across the phase transition in Proposition 4.13. The discontinuity is not quantised, with the leading-order corrections are quadratic in the spin-non-conserving terms. Using continuity arguments and exact results in the absence of Rashba interaction, we conclude that the spin conductivity is not universally quantised for such models. A key challenge in proving Proposition 4.13 is the presence of an arbitrarily small spectral gap, which is in particular smaller than the size of the spin-non conserving terms; in other words, $\|H^{\mathrm{snc}}\|\|(H^{\mathrm{sc}}-\mu\mathbb{1})^{-1}\|$ is not small, making perturbative arguments useless.
@@
Structure of the paper. In Section 2 we establish the framework for describing non-interacting fermions on the honeycomb structure and introduce the $\mathrm{AII}^{\hexagon}$ class. In Section 3 we introduce the spin conductivity via the Kubo formula, and prove that for insulators in the $\mathrm{AII}^{\hexagon}$ class they are well-defined and coincide for both proper and conventional spin current operator (part $(i)$). We then analyse models in the $\mathrm{AII}^{\hexagon}$ class for which the spin is almost conserved and show that that the deviation of the spin conductivity from the quantised value is at worst quadratic in the spin-non-conserving terms (part $(ii)$). In Section 4 we introduce the extended Kane–Mele model, study its phase diagram, and prove that the deviations from the quantised value are quadratic in the spin-non-conserving terms (part $(iii)$). The latter result is based on an imaginary-time representation of the spin conductivity, which is proven in Appendix A via the Wick rotation.
Acknowledgements. We are deeply grateful to M.  Porta for inspiring discussions on transport for electron systems. This work would not have been possible without his constant support. G. M. is grateful to G. Panati for introducing her to this research topic and for stimulating discussions. We acknowledge the financial support from the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (ERC Starting Grant MaMBoQ, no. 802901). The work of L. F. has been supported by the German Research Foundation (DFG) under Germany’s Excellence Strategy - GZ 2047/1, Project-ID 390685813, and under SFB 1060 - Project-ID 211504053. G. M. acknowledges financial support from the Independent Research Fund Denmark–Natural Sciences, grant DFF–10.46540/2032-00005B and from the European Research Council through the ERC CoG UniCoSM, grant agreement n.724939.
2. The $\mathrm{AII}$ class on the honeycomb structure
In this section, we introduce the Hilbert space to describe electrons on the honeycomb structure, discuss the relevant symmetries and introduce the $\mathrm{AII}^{\hexagon}$ models, that is, models in the $\mathrm{AII}$ symmetry class which furthermore possess spatial symmetries compatible with the honeycomb structure. Because of the latter symmetries, such models have desirable properties for studying spin transport, see Section 3. Our notation (as well as Figure 1) is largely borrowed from [MaPaTa, Appendix A].
2.1. Honeycomb structure and Hilbert space
@@
[…]
@@
[…]
@@
We denote the sub-algebra of short-range operators by
[…]
@@
$[A,T_{\gamma}]=0\quad\forall\gamma\in\Gamma\;.$
We denote the sub-algebras of $\Gamma$-periodic bounded and short-range operators respectively by
@@
2.2. The $A\mathrm{II}^{\hexagon}$ class
Besides translations, introduced in (2.7), additional symmetries play a crucial role in understanding the properties of the models we consider. Above all, one should mention the anti-unitary symmetries, upon which the Altland–Cartan–Zirnbauer classification into ten symmetry classes is based [AZ, Kit, RSFL]. Among these classes, the so-called $\mathrm{AII}$ symmetry class is particularly pertinent to our investigation, as the models we consider exhibit time-reversal symmetry but lack particle-hole symmetry. Spatial symmetries are an additional important ingredient for describing the spin-transport properties of the systems, thus necessitating the introduction of the $\mathrm{AII}^{\hexagon}$ (sub-)class in Definition 2.4 below. Note, in passing, that the general importance of spatial symmetries in the understanding of gapped phases of matter was pointed out in [Fu, ShiSaGo].
Before describing the relevant symmetries, let us fix the notation. We let
@@ 2.2. The $A\mathrm{II}^{\hexagon}$ class
$S_{i}:=\frac{1}{2}\mathbb{1}_{\ell^{2}({\mathcal{C}})}\otimes\sigma_{i}\;.$ (2.11)
We also set $\sigma:=(\sigma_{x},\sigma_{y},\sigma_{z})$ and $S:=(S_{x},S_{y},S_{z})$.
Spatial symmetries. We introduce rotation and inversion symmetries as follows:
  1. (i)
    Let $\mathrm{R}_{2\pi/3}$ be the counterclockwise rotation of $32\pi/$ 3 in the plane:
@@
    Since $\mathrm{R}_{2\pi/3}{\mathcal{C}}={\mathcal{C}}$, the restriction of the above map to ${\mathcal{C}}$ is still bijective and will be denoted by the same symbol. The corresponding $32\pi/$ 3-rotation operator on $\mathcal{H}$ is defined as
@@
    An operator $A$ on $\mathcal{H}$ is called $32\pi/$ 3-rotationally symmetric if and only if
    $[A,{R}_{2\pi/3}]=0\;.$
  2. (ii)
    Let $\mathrm{R}_{\mathrm{v}}$ be the vertical reflection in the plane:
    $\left(\mathrm{R}_{\mathrm{v}}\psi\right)(x_{1},x_{2}):=(x_{1},-x_{2})\quad\forall\,(x_{1},x_{2})\in\mathbb{R}^{2}.$
    As $\mathrm{R}_{\mathrm{v}}{\mathcal{C}}={\mathcal{C}}$, the restriction of this function to ${\mathcal{C}}$ is still bijective and will be denoted by the same symbol. We then introduce the spatial-vertical and spin-horizontal reflection (or for brevity $\mathrm{v\otimes h}$-reflection) operator on $\mathcal{H}$, defined as
    $\left({R}_{\mathrm{v\otimes h}}\psi\right)(x):=\mathrm{e}^{-\mathrm{i}\pi S_{y}}\psi(\mathrm{R}^{-1}_{\mathrm{v}}x)=-\mathrm{i}\sigma_{y}\psi(\mathrm{R}^{-1}_{\mathrm{v}}x),\quad\forall\,\psi\in\mathcal{H}.$ (2.13)
    An operator $A$ on $\mathcal{H}$ is called $\mathrm{v\otimes h}$-symmetric if and only if
    $[A,{R}_{\mathrm{v\otimes h}}]=0\;.$
@@
We briefly clarify the name of the reflection symmetry ${R}_{\mathrm{v\otimes h}}$. First of all, we embed the configuration space ${\mathcal{C}}$ in $\mathbb{R}^{3}$ by the canonical injection $(x_{1},x_{2})\mapsto(x_{1},x_{2},0)$ and denote by $\mathrm{R}_{\theta,\hat{n}}$ a counter-clockwise rotation of $\mathbb{R}^{3}$ by an angle $\theta$ around the axis identified by the unit vector $\hat{n}$. This is useful, because a vertical or horizontal reflection on ${\mathcal{C}}$ is then identified with a three-dimensional $\pi$ rotation around the $e_{1}=(1,0,0)$ or $e_{2}=(0,1,0)$ axis respectively. The action of such three-dimensional rotations is extended to $\mathcal{H}$ by setting
@@
whenever $x,\mathrm{R}^{-1}_{\theta,\hat{n}}x\in{\mathcal{C}}$. In particular, one has that ${R}_{2\pi/3}=R_{2\pi/3,e_{3}}$ where $e_{3}=(0,0,1)$. On the other hand, the transformation ${R}_{\mathrm{v\otimes h}}$ performs a vertical reflection in space and a horizontal reflection in spin.
Anti-unitary symmetries. We introduce time-reversal and particle-hole symmetries:
  1. (iii)
    The time-reversal operator is defined as
    $\left({\Theta}\psi\right)(x):=\mathrm{e}^{\mathrm{i}\pi S_{y}}\left(K\psi\right)(x)=\left(\mathbb{1}\otimes\mathrm{i}\sigma_{y}\right)\overline{\psi(x)},\quad\forall\,\psi\in\mathcal{H}$
    where $K$ is the complex conjugation and $\Theta$ is such that $\Theta^{2}=-\mathbb{1}_{\mathcal{H}}$. An operator $A$ on $\mathcal{H}$ is called time-reversal symmetric if and only if
    $[A,\Theta]=0\;.$
  2. (iv)
    The particle-hole operator is defined as [HK]
    $\left(\Xi\psi\right)(x):=(\mathbb{1}\otimes\sigma_{x})\overline{\psi(x)}\quad\forall\,\psi\in\mathcal{H},$
    which is such that $\Xi^{2}=\mathbb{1}_{\mathcal{H}}$. An operator $A$ on $\mathcal{H}$ is said to be particle-hole symmetric if and only if
    $\{A,\Xi\}=0\;.$
Remark \@upn2.3.
Operators acting on $\mathcal{H}$ that are time-reversal symmetric with $\Theta^{2}=-\mathbb{1}$ but not particle-hole symmetric belong to the Altland–Cartan–Zirnbauer class $\mathrm{AII}$, see, e. g. [HK].
We can finally introduce the $\mathrm{AII^{\hexagon}}$ class.
Definition \@upn2.4 ($\mathrm{AII^{\hexagon}}$ class).
The $\mathrm{AII^{\hexagon}}$ class consists of all the operators in $\mathcal{P}(\mathcal{H})$ which are $32\pi/$ 3-rotationally, $\mathrm{v\otimes h}$- and time-reversal symmetric, but not particle-hole symmetric, that is,
$\mathrm{AII}^{\hexagon}:=\big{\{}H\in\mathcal{P}(\mathcal{H})\,|\,[H,R_{2\pi/3}]=[H,{R}_{\mathrm{v\otimes h}}]=[H,\Theta]=0\,,\{H,\Xi\}\neq 0\big{\}}\;.$
The goal of the paper is to show that spin transport can be put on solid mathematical grounds within the $\mathrm{AII}^{\hexagon}$ class, see Section 3, while also establishing that the spin conductivity lacks universality property, see Section 4.2. The Kane–Mele model [KM1, KM2] is a paradigmatic example of a Hamiltonian in the $\mathrm{AII}^{\hexagon}$ class. In Section 4 we will consider an extension of the Kane–Mele model and show that it belongs to the $\mathrm{AII}^{\hexagon}$ class, see Lemma 4.3.
3. Spin transport in $\mathrm{AII}^{\hexagon}$

### v2

# Spin transport and lack of quantisation for time-reversal symmetric insulators on the honeycomb structure

@@ Abstract.
We investigate spin transport in a class of time-reversal symmetric insulators on the honeycomb structure, the Kane–Mele model being an emblematic example in this class. We derive the spin conductivity by the linear response à la Kubo and show that it is well-defined and independent of the choice of the spin current. For models that do not conserve the spin, we demonstrate that the deviation of the spin conductivity from the quantised value is, at worst, quadratic in the spin-non-conserving terms, thus improving previous results. Additionally, we show that the leading-order corrections are actually quadratic for some models in the class, demonstrating that the spin conductivity is not universally quantised. Consequently, our results show that, in general, there is no direct connection between the spin conductivity and the Fu–Kane–Mele index.
@@ 1. Introduction
For spin-conserving systems, the spin Hall conductivity is connected with the Fu–Kane–Mele index, as is the case for the charge Hall conductivity with the Chern number: in particular, the spin Hall conductivity is quantised in units of $e/h$, the quantisation integer modulo two being the Fu–Kane–Mele index [BHZ2, HK]. However, in spin-non-conserving systems, it remains both experimentally and theoretically unclear whether any spin response function can be associated with the Fu–Kane–Mele index. On the one hand, experimental measurements are imprecise, likely because of the presence of magnetic impurities that partially break time-reversal symmetry [Mol1, Mol2]. On the other hand, theoretical problems already arise in identifying a suitable spin-current operator, the system lacking the associated conserved quantity [SZXN]. In this regard, two operators serve as natural generalizations of the spin current for spin-conserving systems: the conventional spin current, defined as the product of the charge current with the spin operator, and the proper spin current introduced in [SZXN, ZWSXN], which is tied with a continuity equation and with the Onsager relations.
@@
In this work, using the Kubo formula, we continue the study of the spin conductivity initiated in [MaPaTa, MaPaTe, MaMo], by specialising to short-range, translation and $2\pi/3$-rotation symmetric models on the two-dimensional honeycomb structure. The presence of spatial symmetries is crucial in our analysis to remove the ill-posedness problems and the ambiguities related to the choice of the spin current. Unlike [MaPaTe], in which the NEASS approach is used, here we follow Kubo’s approach to transport coefficients. Moreover, by further exploiting spatial inversion symmetries of the honeycomb structure, we prove that the spin conductivity is an antisymmetric tensor. This property makes the spin conductivity invariant under rotations, and thus establishes that it measures an intrinsic transverse response, independent of the orientation of the laboratory [SVWBJ].
In addition to reviewing key well-posedness findings, we prove two novel results for short-range, translation- and $2\pi/3$-rotation symmetric insulators that are time-reversal symmetric and that almost conserve the spin. First, we show that the deviation of the spin conductivity from the quantised value (which is observed in spin-conserving systems) is, at worst, quadratic in the spin-non-conserving terms, improving previous results [Sch, MaPaTe]. Note that in theoretical units, which we henceforth adopt, quantisation of the spin conductivity means valued in $\frac{1}{2\pi}\mathbb{Z}$. Additionally, we show that in certain models the leading-order corrections to the quantised value are indeed quadratic, implying that spin conductivity is not universally quantised. As a result, our findings demonstrate that no general connection exists between spin conductivity and the Fu–Kane–Mele index when implementing the linear response by Kubo’s formula. To the best of our knowledge, this is the first rigorous result in this direction.
Let us now state these results more precisely. We refer to insulators as pairs $(H,\mu)$ where $H$ is a Hamiltonian with a spectral gap, and where $\mu\in\mathbb{R}$ is in the spectral gap. An insulator $(H,\mu)$ is short-range, translation- and $2\pi/3$-rotation symmetric if $H$ is such, see (2.13) and Section 2 for details.
Let us denote by $S_{z}=\frac{1}{2}\sigma_{z}$ the spin operator in the $z$ direction. To quantify spin conservation, we follow [Sch, Eq. (1)] and split any Hamilton operator $H$ as follows
@@
    For any short-range, translation- and $2\pi/3$-rotation symmetric insulator $(H,\mu)$, the spin conductivity tensor, which we denote by $\sigma^{s}(H,\mu)=\big(\sigma^{s}_{ij}(H,\mu)\big)_{i,j=1,2}$, is well-defined and independent of the choice of the spin current operator. Moreover, if $H$ is (spatially) inversion symmetric it holds true that $\sigma^{s}_{ij}(H,\mu)=-\sigma^{s}_{ji}(H,\mu)$, for any $i,j=1,2$.
@@
[…]
@@
[…]
Each item of Theorem 1.1 is elaborated in different sections of the manuscript: part $(i)$ is addressed in Propositions 3.13 and 3.15, part $(ii)$ in Theorem 3.22, and finally part $(iii)$ in Theorem 4.10. As previously mentioned, the proof of $(i)$ follows the ideas of [MaPaTa, MaPaTe], but relies on Kubo’s approach to the computation of transport coefficients. Our analysis shows that the derivation of the spin conductivity by Kubo’s formula agrees with the one obtained via NEASS, see Remark 3.18. The proof of $(ii)$ involves more refined algebraic manipulations than in previous work, allowing us to extract commutators between operators that nearly commute with the spin operator, thus providing improved estimates. Iterating these manipulations does not lead to further cancellations or sharper bounds; yet, $(ii)$ does not rule out the possibility that the corrections could be universally of higher order. This is however not the case due to $(iii)$. To prove $(iii)$, we explicitly construct an extension of the Kane–Mele model, introduced in [KM1, KM2], with next-to-nearest-neighbour Rashba interaction (see (4.1)) and analyse the phase diagram for suitable parameter choices. In Proposition 4.9, we show that, like in the Kane–Mele model, this diagram features three insulating phases separated by two semi-metallic phases (i.e., a metallic phase with a point-like Fermi surface). Using the “imaginary-time” representation of the Kubo formula for the spin conductivity, established via a Wick rotation (as in [GJMP, GMP2]) and detailed in Theorem 4.15, we compute the discontinuity of the spin conductivity across the phase transition in Proposition 4.13. The discontinuity is not quantised, with the leading-order corrections are quadratic in the spin-non-conserving terms. Using continuity arguments and exact results in the absence of Rashba interaction, we conclude that the spin conductivity is not universally quantised for such models. A key challenge in proving Proposition 4.13 is the presence of an arbitrarily small spectral gap, which is in particular smaller than the size of the spin-non conserving terms; in other words, $\|H^{\mathrm{snc}}\|\|(H^{\mathrm{sc}}-\mu\mathbb{1})^{-1}\|$ is not small, making perturbative arguments useless.
@@
Structure of the paper. In Section 2 we establish the framework for describing non-interacting fermions on the honeycomb structure and introduce the algebra of short-range, translation and $2\pi/3$-rotation symmetric linear operators $\mathcal{P}_{0}^{\hexagon}$. In Section 3 we introduce the spin conductivity via the Kubo formula, and prove that for insulators in $\mathcal{P}_{0}^{\hexagon}$ they are well-defined and coincide for both proper and conventional spin current operator (Theorem 1.1(i)). We then analyse time-reversal symmetric models in $\mathcal{P}_{0}^{\hexagon}$ for which the spin is almost conserved and show that the deviation of the spin conductivity from the quantised value is at worst quadratic in the spin-non-conserving terms (Theorem 1.1(ii)). In Section 4 we introduce the extended Kane–Mele model, study its phase diagram, and prove that the deviations from the quantised value are quadratic in the spin-non-conserving terms (Theorem 1.1(iii)). The latter result is based on an imaginary-time representation of the spin conductivity, which is proven in Appendix A via the Wick rotation.
Acknowledgements. We are deeply grateful to M.  Porta for inspiring discussions on transport for electron systems and for his constant support, without which this work would not have been possible. G. M. is grateful to G. Panati for introducing her to this research topic and for stimulating discussions. We also acknowledge G.M. Graf and D. Monaco for valuable and stimulating discussions. We acknowledge the financial support from the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (ERC Starting Grant MaMBoQ, no. 802901). The work of L. F. has been supported by the German Research Foundation (DFG) under Germany’s Excellence Strategy - GZ 2047/1, Project-ID 390685813, and under SFB 1060 - Project-ID 211504053. G. M. acknowledges financial support from the Independent Research Fund Denmark–Natural Sciences, grant DFF–10.46540/2032-00005B and from the European Research Council through the ERC CoG UniCoSM, grant agreement n.724939. We finally thank the anonymous referee for useful comments.
2. The honeycomb structure
In this section, we introduce the Hilbert space $\mathcal{H}$ to describe electrons on the honeycomb structure, the relevant symmetries and the subalgebra $\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$ of short-range translation- and $2\pi/3$-rotation symmetric operators. Models in $\mathcal{P}_{0}^{\hexagon}(\mathcal{H})$ have desirable properties for studying spin transport, see Section 3. Our notation (as well as Figure 1) is largely borrowed from [MaPaTa, Appendix A].
@@ 2. The honeycomb structure
We denote the algebra of linear bounded operators acting on $\mathcal{H}$ by $\mathcal{B}(\mathcal{H})$ and by $\|\cdot\|$ the operator norm. For any operator $A\in\mathcal{B}(\mathcal{H})$, we denote by ${[A]}_{{\mathcal{D}}_{i}}:={\mathcal{D}}_{i}\,A\,{\mathcal{D}}_{i}^{-1}$ the corresponding operator acting on $\ell^{2}(\mathbb{Z}^{2})\otimes\mathbb{C}^{4}$. We also adopt the following notation: for every $n,m\in\mathbb{Z}^{2}$
@@
where $\delta_{n}$ is defined as usual ${(\delta_{n})}_{m}=\delta_{n,m}$ and $\left\langle\,\cdot\,$$\cdot\,\right\rangle$ is the standard scalar product on $\ell^{2}(\mathbb{Z}^{2})$. Obviously, $A$ is characterized by the family of matrices $\{A^{(i)}_{n,m}\}_{n,m\in\mathbb{Z}^{2}}\subset\operatorname{End}\big(\mathbb{C}^{4}\big)$. Unless specified, we will consider the $i=3$ dimerization and, by abuse of notation, refer simply to $\{A_{n,m}\}_{n,m\in\mathbb{Z}^{2}}$. For such matrices we use the notation $\left\lvert A_{n,m}\right\rvert$ for the matrix norm.
@@
We denote the subalgebra of short-range operators by
$\mathcal{B}_{0}(\mathcal{H}):=\{A\in\mathcal{B}(\mathcal{H})\,|\,\text{$A$ is short-range}\}\;.$
@@
$A,T_{\gamma}]=0\quad\forall\gamma\in\Gamma\;.$
We denote the subalgebras of $\Gamma$-periodic bounded and short-range operators respectively by
@@
Besides translations, other symmetries will play an important role in the rest of this work. These symmetries act non-trivially on the spin degrees of freedom. We let
@@
$S_{i}:=\frac{1}{2}\mathbb{1}_{\ell^{2}({\mathcal{C}})}\otimes\sigma_{i}\;,$ (2.11)
and write $\sigma:=(\sigma_{x},\sigma_{y},\sigma_{z})$ and $S:=(S_{x},S_{y},S_{z})$. We also let $\mathrm{R}_{2\pi/3}$ denote the counterclockwise rotation of $2\pi/3$ in the plane:
@@
Since $\mathrm{R}_{2\pi/3}{\mathcal{C}}={\mathcal{C}}$, the restriction of the above map to $\mathcal{C}$} is still bijective and will be denoted by the same symbol. We introduce the $2\pi/3$-rotation operator on $\mathcal{H}$
@@
Note that we embed the configuration space $\mathcal{C}$} in $\mathbb{R}^{3}$ by the canonical injection $(x_{1},x_{2})\mapsto(x_{1},x_{2},0)$ and denote by $\mathrm{R}_{\theta,\hat{n}}$ a counter-clockwise rotation of $\mathbb{R}^{3}$ by an angle $\theta$ around the axis identified by the unit vector $\hat{n}$. The action of such three-dimensional rotations is defined on $\mathcal{H}$ by setting
@@
whenever $x,\mathrm{R}^{-1}_{\theta,\hat{n}}x\in{\mathcal{C}}$. In particular, one has that ${R}_{2\pi/3}=R_{2\pi/3,e_{3}}$ where $e_{3}=(0,0,1)$.
A linear operator $A$ on $\mathcal{H}$ is called $2\pi/3$-rotation symmetric if and only if
$A,{R}_{2\pi/3}]=0\;.$
Together with translations, rotational symmetry plays a crucial role in understanding the spin-transport properties of the models we consider. We therefore introduce the subalgebra of short-range $\Gamma$-periodic, $2\pi/3$-rotation symmetric operators
$\begin{split}\mathcal{P}^{\hexagon}_{0}(\mathcal{H})&:=\{A\in\mathcal{P}_{0}(\mathcal{H})\,|\,[A,{R}_{2\pi/3}]=0\}\;,\end{split}$ (2.13)
which will be our main focus throughout the rest of this work.
Other important symmetries of the honeycomb structure are vertical and horizontal spatial inversions. Let $a$ be either 0 or $1$. We say that $\Pi_{a}$ is an inversion if and only if $\Pi_{a}$ is an operator on $\mathcal{H}$ such that, possibly after an overall translation (2) (2) (2)In fact, for $a=0$, that is, for the horizontal inversion, one needs to translate by (1/2,0)(-1/2,0)., $\Pi_{a}^{2}=\mathbb{1}$,
$\Pi_{a}\,X_{i}\,\Pi_{a}^{-1}=(-)^{i+a}X_{i}\quad\text{ and }\quad\Pi_{a}\,S_{z}\,\Pi_{a}^{-1}=-S_{z}.$ (2.14)
The inversion $\Pi_{0}$ is referred to as horizontal, while $\Pi_{1}$ is referred to as vertical. An operator $A$ on $\mathcal{H}$ is called inversion symmetric if and only if
$A,\Pi_{a}]=0\qquad\text{for some $a\in\{0,1\}$}.$
Spatial inversion symmetries will be used to prove that the spin conductivity is an antisymmetric tensor, see Proposition 3.15, and will be also present in the extended Kane–Mele model discussed in Section 4.1.
Finally, we introduce time-reversal symmetry on $\mathcal{H}$. The time-reversal operator is an anti-unitary operator that flips the spin and acts trivially on the spatial degrees of freedom. For spin $1/2$ fermions on the honeycomb structure it is given by
$(\Theta\psi)(x)=\mathrm{e}^{\mathrm{i}\pi S_{y}}\overline{\psi(x)}\;,\qquad\forall\psi\in\mathcal{H}\,$, (2.15)
and satisfies $\Theta^{2}=-\mathbb{1}_{\mathcal{H}}$. An operator $A$ on $\mathcal{H}$ is called time-reversal symmetric if and only if
$A,\Theta]=0\;.$
3. Spin transport